OpenStudy (anonymous):

solve for x. log[4](x+4)+log[4](x+16)=3 a. 16 b. 0 or -20 c. 0 d. -20 e. none of the above

OpenStudy (radar):

c, 1+2=3

OpenStudy (radar):

\[\log _{4}(0+4)=1\]\[\log _{4}(0+16)=2\]

OpenStudy (radar):

Need to check that -20 out also

jhonyy9 (jhonyy9):

nice

OpenStudy (radar):

Need to check that -20 out also\[\log _{4}(-20+4)=\log _{4}-16\] don't think that is a 2. Anyone want to comment.

OpenStudy (radar):

\[4^{x}=-16\] ?????

jhonyy9 (jhonyy9):

4*4*(-1)

jhonyy9 (jhonyy9):

is only 0

jhonyy9 (jhonyy9):

or just

OpenStudy (radar):

O.K Thanks

jhonyy9 (jhonyy9):

nice

OpenStudy (radar):

I guess that wraps it up Yardbird84

OpenStudy (anonymous):

thank you for your help